# Two-dimensional fluids via matrix hydrodynamics

### v2

@@ 6. A conjecture on the long-time behavior and integrability
Shnirelman [35] conjectured that $\Omega_{+}(\omega_{0})$ is compact in $L^{2}$. This conjecture can be understood as follows. $\Omega_{+}(\omega_{0})$ is indeed compact in the weak $L^{2}$ topology, so the $L^{2}$ norm (enstrophy) attains a minimum on it, denoted by $\hat{\omega}$. We observe that $\overline{\{S_{t}(\hat{\omega})\mid t\geq 0\}}^{*}\subset\Omega_{+}(\omega_{0})$ and also the $L^{2}$ norm is constant on $\overline{\{S_{t}(\hat{\omega})\mid t\geq 0\}}^{*}$. Therefore, any weakly convergent sequence in $\overline{\{S_{t}(\hat{\omega})\mid t\geq 0\}}^{*}$ is also strongly convergent in $L^{2}$. Consequently, the conjecture of Shnirelman is that no further mixing (loss of enstrophy) can occur for initial data in $\Omega_{+}(\omega_{0})$.
The conjecture does not imply that steady state is reached (e.g. a minimal flow). Indeed, unsteadiness can persist indefinitely, without further mixing. From batches of numerical experiments [30], we conjecture that persistent unsteadiness in $\Omega_{+}(\omega_{0})$ is connected to integrable configuration of “vortex blob dynamics”, where the centers of mass of $n$ fixed-shape blobs interact (cf. Fig. 5(a)). Depending on the domain’s geometry, there is a threshold $n$ for which the dynamics is always integrable [31]. The underlying mechanism is then that vortex mixing continues until a large-scale almost integrable configuration of vortex blobs is reached. Being quasi-periodic, it then acts as a barrier for further mixing, trapping the vorticity in persistent unsteadiness. On $\mathbb{S}^{2}$, the thresholds are $n=3$ for generic initial data, and $n=4$ for vanishing angular momentum. Long-time numerical simulations with Zeitlin’s model align with this conjecture: for generic (smooth) initial data they typically settle on $n=3$ blobs (e.g. Fig. 1), and for vanishing momentum on $n=4$ blobs; see [30, Sec. 4] for details.
@@ Appendix A Proof of Theorem 4
Let $T_{N}$ denotes the projection of a function from $L^{2}(\mathbb{S}^{2})$ to $\mathfrak{su}(N)$. In the following calculations, we mainly follow the approach of Gallagher [13] and the results by Flandoli, Pappalettera, and Viviani [12, Appendix  A]. We begin with the following result (with the convention that $\lVert\cdot\rVert_{s}$ denotes the $H^{s}(\mathbb{S}^{2})$ Sobolev norm and $\lVert\cdot\rVert_{0}$ when applied to matrices is the scaled Frobenius norm).
@@
Let $f\in L^{2}(\mathbb{S}^{2})$ and $g\in H^{3+\varepsilon}(\mathbb{S}^{2})$, we have that for any $\varepsilon>0$, $30<\alpha<2/$ 3, and $50<\beta<2/$ 5 there exist positive constants $C(\alpha)$, $C(\beta)$, $C(\alpha,\epsilon)$, and $C(\beta,\epsilon)$ such that
@@
Let $N$ be an even integer, then we want to estimate the difference:
@@
where $10<\alpha,\beta<$ 1 have to be determined. Notice that the sum
@@
Step 1. For the first term, we have to divide sum in two cases: $l_{2}<l_{1}\land N^{\beta}$ and $l_{1}\leq l_{2}\land N^{\alpha}$. Using [12, Prop. 14 (1)], we get
@@
Step 2. For the second term, we proceed analogously, using the result in [12, Prop.14 (2)]. Since, $\beta<\alpha$, we always have $l_{2}<l_{1}$. Therefore,
@@
for some $\varepsilon>0$, $30<\alpha<2/$ 3 and $50<\beta<2/$ 5.
@@
Repeating the same calculations of Lemma 13, dividing in the two cases $l_{1}<l_{2}$ and $l_{2}\leq l_{1}$, and using the result in [12, Lemma 12], we get that:
@@
Let $\omega=\omega(t,x)$ and $W=W(t)$ be the solutions to the Euler equations (5) and Euler–Zeitlin equations (7), with initial data $\omega(0)\in H^{3+\varepsilon}(\mathbb{S}^{2})$ for some $\varepsilon>0$ and $W(0)\in\mathfrak{su}(N)$. Let $\omega_{N}(t):=T_{N}^{*}W(t)$. Then, for any $t\geq 0$, and any $\varepsilon>0$, $30<\alpha<2/$ 3 and $50<\beta<2/$ 5, there exist positive constants $C(\alpha),C(\beta),C(\alpha,\varepsilon),C(\beta,\varepsilon)$ such that
@@
[…]

### v3

@@ 6. A conjecture on the long-time behavior and integrability
Shnirelman [50] conjectured that $\Omega_{+}(\omega_{0})$ is compact in $L^{2}$. This conjecture can be understood as follows. $\Omega_{+}(\omega_{0})$ is indeed compact in the weak $L^{2}$ topology, so the $L^{2}$ norm (enstrophy) attains a minimum $\hat{\omega}$ on it. We observe that $\overline{\{S_{t}(\hat{\omega})\mid t\geq 0\}}^{*}\subset\Omega_{+}(\omega_{0})$ and also the $L^{2}$ norm is constant on $\overline{\{S_{t}(\hat{\omega})\mid t\geq 0\}}^{*}$. Therefore, any weakly convergent sequence in $\overline{\{S_{t}(\hat{\omega})\mid t\geq 0\}}^{*}$ is also strongly convergent in $L^{2}$. Consequently, the conjecture of Shnirelman is that no further mixing (loss of enstrophy) can occur for initial data in $\Omega_{+}(\omega_{0})$.
The conjecture does not imply that steady state is reached (e.g. a minimal flow). Indeed, unsteadiness can persist indefinitely, without further mixing. From batches of numerical experiments [45], we conjecture that persistent unsteadiness in $\Omega_{+}(\omega_{0})$ is connected to integrable configuration of “vortex blob dynamics”, where the centers of mass of $n$ fixed-shape blobs interact (cf. Fig. 5(a)). Depending on the domain’s geometry, there is a threshold $n$ for which the dynamics is always integrable [46]. The underlying mechanism is then that vortex mixing continues until a large-scale almost integrable configuration of vortex blobs is reached. Being quasi-periodic, it then acts as a barrier for further mixing, trapping the vorticity in persistent unsteadiness. On $\mathbb{S}^{2}$, the thresholds are $n=3$ for generic initial data, and $n=4$ for vanishing angular momentum. Long-time numerical simulations with Zeitlin’s model align with this conjecture: for generic (smooth) initial data they typically settle on $n=3$ blobs (e.g. Fig. 1), and for vanishing momentum on $n=4$ blobs; see [45, Sec. 4] for details.
@@ Appendix A Proof of Theorem 4
Let $T_{N}$ denotes the projection of a function from $L^{2}(\mathbb{S}^{2})$ to $\mathfrak{su}(N)$. In the following calculations, we mainly follow the approach of Gallagher [23] and the results by Flandoli, Pappalettera, and Viviani [22, Appendix  A]. The basic strategy is to show that the Zeitlin flow $W(t)\in\mathfrak{su}(N)$ lifted to $L^{2}$ via $\omega_{N}(t)\coloneqq T_{N}^{*}W(t)$ fulfills a differential inequality of the form
$\frac{d}{dt}\lVert\omega_{N}(t)-\omega(t)\rVert\leq a_{N}+b_{N}\lVert\omega_{N}(t)-\omega(t)\rVert$
such that $a_{N}\to 0$ and $b_{N}$ is bounded as $N\to\infty$ and $0\leq t\leq t_{\it max}<\infty$. Convergence then follow from the Grönwall lemma if $\lim_{N\to\infty}\lVert\omega_{N}(0)-\omega(0)\rVert=0$. To achieve the differential inequality we need to understand how the matrix commutator $\cdot$$\cdot$] interacts with the Poisson bracket $\{\cdot,\cdot\}$. This is done by studying the relation between the corresponding Lie algebra structure constants relative to the spherical harmonics basis (which is available for both functions and matrices).
We begin with the following result (with the convention that $\lVert\cdot\rVert_{s}$ denotes the $H^{s}(\mathbb{S}^{2})$ Sobolev norm, $\lVert\cdot\rVert_{0}$ applied to matrices is the scaled Frobenius norm, and $\cdot\,,\cdot]_{N}\coloneqq[\cdot\,,\cdot]/\hbar_{N}$).
@@
Let $f\in L^{2}(\mathbb{S}^{2})$ and $g\in H^{3+\varepsilon}(\mathbb{S}^{2})$, we have that for any $\varepsilon>0$, $0<\alpha<2/3$, and $0<\beta<2/5$ there exist positive constants $C(\alpha)$, $C(\beta)$, $C(\alpha,\epsilon)$, and $C(\beta,\epsilon)$ such that
@@
Let $C^{(N)l_{3}m_{3}}_{l_{1}m_{1},l_{2}m_{2}}$ and $C^{l_{3}m_{3}}_{l_{1}m_{1},l_{2}m_{2}}$ denote the structure constants of $\cdot,\cdot]_{N}$ and $\{\cdot,\cdot\}$ relative to the spherical harmonics basis (see Hoppe [27] and Yoshida [58]), and let $N$ be an even integer. We want to estimate the difference
@@
where $0<\alpha,\beta<1$ have to be determined. Notice that the sum
@@
Step 1. For the first term, we have to divide sum in two cases: $l_{2}<l_{1}\land N^{\beta}$ and $l_{1}\leq l_{2}\land N^{\alpha}$. Using [22, prop. 14 (1)], we get
@@
Step 2. For the second term, we proceed analogously, using the result in [22, prop. 14 (2)]. Since, $\beta<\alpha$, we always have $l_{2}<l_{1}$. Therefore,
@@
for some $\varepsilon>0$, $0<\alpha<2/3$ and $0<\beta<2/5$.
@@
Repeating the same calculations of Lemma 16, dividing in the two cases $l_{1}<l_{2}$ and $l_{2}\leq l_{1}$, and using the result in [22, Lemma 12], we get that:
@@
Let $\omega=\omega(t,x)$ and $W=W(t)$ be the solutions to the Euler equations (6) and Euler–Zeitlin equations (9), with initial data $\omega(0)\in H^{3+\varepsilon}(\mathbb{S}^{2})$ for some $\varepsilon>0$ and $W(0)\in\mathfrak{su}(N)$. Let $\omega_{N}(t):=T_{N}^{*}W(t)$. Then, for any $t\geq 0$, and any $\varepsilon>0$, $0<\alpha<2/3$ and $0<\beta<2/5$, there exist positive constants $C(\alpha),C(\beta),C(\alpha,\varepsilon),C(\beta,\varepsilon)$ such that
@@
$\|\omega(0)\|_{0}\int_{0}^{t}\dfrac{C(\alpha)}{N^{2(1-3/2\alpha)}}\|\omega(s)\|_{1}ds+\dfrac{t\,C(\beta)}{N^{2(1-5/2\beta)}}\|\omega(0)\|_{0}^{2}+$
@@
Appendix B Relation to Berezin-Toeplitz quantization
Since Dirac [15] outlined quantization theory, it has ramified into different mathematical developments. In this paper, we obtain quantization via unitary representation theory, which perhaps is the most direct route when applicable. Non-commutative geometry (cf. [12]) offers a broader viewpoint, and geometric quantization (cf.  [52]) yet another. In general, the various frameworks are non-equivalent and applicable under different conditions. But for the sphere they coincide; it is somehow the archetype of quantization (in non-commutative geometry the “fuzzy sphere”, and in geometric quantization the tautological complex line bundle over the Riemann sphere $\mathbb{C}P^{1}$, as we shall see).
Theorem 3 above was stated and proved by Charles and Polterovich [8] in the framework of Berezin-Toeplitz quantization (which is a type of geometric quantization, cf. [4, 40]). The purpose of this section is to show how quantization of the sphere via representations, as in section 3.1 above, corresponds to Berezin-Toeplitz quantization of the Riemann sphere. This result is essentially contained in the original work of Berezin [4], and fully explained in the monograph by Le Floch [40], to which we refer for details.
Let $M$ be a compact Kähler manifold. The objective is to construct a mapping (the Berezin-Toeplitz quantization) between the space of smooth functions $C^{\infty}(M,\mathbb{C})$ and a space of operators on a finite dimensional quantum state space (i.e., the Hilbert space of “wave-functions”). State space is constructed as holomorphic sections of a tensor power of a Hermitian line bundle over $M$. Whereas this construction appears abstruse at first, it is straightforward in the simplest setting $M=\mathbb{C}P^{1}$. As we shall see, this corresponds to quantization on $\mathbb{S}^{2}$ as presented in section 3.1.
Figure 6. Illustration of the tautological line bundle $\mathcal{O}(-1)$. An element $([\mathbf{z}],v)\in\mathcal{O}(-1)$ consist of a one-dimensional complex linear subspace $\mathbf{z}]\in\mathbb{C}P^{1}$ together with a vector in that subspace $v\in[\mathbf{z}]$. Thus, a linear form $\ell\in(\mathbb{C}^{2})^{*}$ naturally pairs with elements of $\mathcal{O}(-1)$.
[…]
There is a natural vector bundle associated with the Riemann sphere: the fiber above $\mathbf{z}]\in\mathbb{C}P^{1}$ is a vector $v\in[\mathbf{z}]$ of the co-set. This is the tautological line bundle, denoted $\mathcal{O}(-1)$ and illustrated in Fig. 6. Since it is a sub-bundle of the trivial vector bundle $\mathbb{C}P^{1}\times\mathbb{C}^{2}$, it inherits the Hermitian inner product of $\mathbb{C}^{2}$. Furthermore, it is a holomorphic line bundle in the sense that it is a holomorphic manifold and the projection $\mathcal{O}(-1)\to\mathbb{C}P^{1}$ is a holomorphic mapping.
Consider now an element $p\in(\mathbb{C}^{2})^{*}$, i.e., a linear function $p\colon\mathbb{C}^{2}\to\mathbb{C}$. As illustrated in Fig. 6, there is a natural bundle pairing between $p$ and a smooth section $s\in C^{\infty}(\mathbb{C}P^{1},\mathcal{O}(-1))$, namely
$s([\mathbf{z}])\mapsto p(s([\mathbf{z}])).$
[…]
$\mathrm{KS}_{1}(f)=f-i\nabla_{X_{f}}$
where $f$ acts by multiplication, $\nabla$ is the connection associated with the Hermitian structure of the line bundle $\mathcal{O}(1)$ (i.e., the Chern connection), and $X_{f}$ is the Hamiltonian vector field for $f$. The Kostant-Souriau operators fulfill the quantization condition [^7] The Hermitian structure is scaled so the associated symplectic structure $\Omega$ fulfills $\int_{\mathbb{C}P^{1}}\Omega=2\pi$.
$\mathrm{KS}_{1}(f),\mathrm{KS}_{1}(g)]=-i\mathrm{KS}_{1}(\{f,g\}).$
But the space $C^{\infty}\big{(}\mathbb{C}P^{1},\mathcal{O}(1)\big{)}$ is inadequate as quantum state space: it is infinite-dimensional and admits far too many self-adjoint operators. In contrast, the space of holomorphic sections of $\mathcal{O}(1)$ is too small (as it is given by $(\mathbb{C}^{2})^{*}$). We need something in-between. The solution is to consider the tensor product $\mathcal{O}(-k)\equiv\mathcal{O}(-1)^{\otimes k}$ and then take holomorphic sections of its dual bundle $\mathcal{O}(k)\equiv\mathcal{O}(-k)^{*}$. Notice that $\mathcal{O}(k)$ is still a line bundle (since the fiber of $\mathcal{O}(-1)$ is one-dimensional), so this construction might appear fruitless. But the point is that it admits more holomorphic sections. Indeed, every homogenous polynomial $p\in\mathbb{C}_{k}[z_{1},z_{2}]$ of degree $k$ gives rise to a $k$-linear form on $\mathcal{O}(-1)$ via
$\mathbf{z}]^{k}\ni(\lambda_{1}\mathbf{z},\ldots,\lambda_{k}\mathbf{z})\mapsto\lambda_{1}\cdots\lambda_{k}p(\mathbf{z}).$
Due to the homogeneity of $p$, this definition is independent of the representative $\mathbf{z}\in[\mathbf{z}]$. It defines a holomorphic section of $\mathcal{O}(k)$ (using that $k$-linear forms on $\mathcal{O}(-1)$ are naturally identified with sections of $(\mathcal{O}(-1)^{\otimes k})^{*}=\mathcal{O}(k)$). Moreover, all holomorphic sections of $\mathcal{O}(k)$ are of this form, so the space of homogenous polynomials $\mathbb{C}_{k}[z_{1},z_{2}]$ is isomorphic to the space of holomorphic sections of $\mathcal{O}(k)$.
With this identification, consider the $L^{2}$ projection
$\Pi\colon L^{2}(\mathbb{C}P^{1},\mathcal{O}(k))\to\mathbb{C}_{k}[z_{1},z_{2}]\qquad\text{(the Szegő projector.)}$
The Berezin-Toeplitz operator for $f\in C^{\infty}(\mathbb{C}P^{1},\mathbb{C})$ is
$\mathrm{BT}_{k}(f)\colon\mathbb{C}_{k}[z_{1},z_{2}]\to\mathbb{C}_{k}[z_{1},z_{2}],\quad\mathrm{BT}_{k}(f)=\Pi\circ f,$
where $f$ acts by multiplication. If $f$ is real-valued, then $\mathrm{BT}_{k}(f)$ is self-adjoint.
Let $\mathrm{KS}_{k}(f)$ denote the generalization of $\mathrm{KS}_{1}(f)$ to the operator
$\mathrm{KS}_{k}(f)=f-\frac{i}{k}\nabla^{k}_{X_{f}}\colon C^{\infty}(\mathbb{C}P^{1},\mathcal{O}(k))\to C^{\infty}(\mathbb{C}P^{1},\mathcal{O}(k)),$
where $\nabla^{k}$ is the connection on $\mathcal{O}(k)$. Notice that the operator $\mathrm{BT}_{k}(f)$ is different from “standard” geometric quantization, which is $\Pi\circ\mathrm{KS}_{k}$ as an operator $\mathbb{C}_{k}[z_{1},z_{2}]\to\mathbb{C}_{k}[z_{1},z_{2}]$. Via Tuynman’s [56] formula $\Pi\circ 2i\nabla^{k}_{X_{f}}\circ\Pi=\Pi\circ(\Delta f)\circ\Pi$ one concludes the relation $\Pi\circ\mathrm{KS}_{k}(f)=\mathrm{BT}_{k}(f)-\frac{1}{2k}\mathrm{BT}_{k}(\Delta f)$.
To obtain the connection to quantization via representation as described above, consider the Berezin-Toeplitz operators associated with the coordinates functions $x_{1},x_{2},x_{3}$ of the sphere (again identified with $\mathbb{C}P^{1}$). In the coordinate chart where $p\in\mathbb{C}_{k}[z_{1},z_{2}]$ is identified with the one-variable polynomial $z_{1}\mapsto p(z_{1},1)$, they are given by
$\mathrm{BT}_{k}(x_{1})=\frac{1}{k+2}\left((1-z_{1}^{2})\frac{d}{dz_{1}}+kz_{1}\right),$
$\mathrm{BT}_{k}(x_{2})=\frac{i}{k+2}\left((1+z_{1}^{2})\frac{d}{dz_{1}}-kz_{1}\right),$
$\mathrm{BT}_{k}(x_{3})=\frac{1}{k+2}\left(2z_{1}\frac{d}{dz_{1}}-k\right)\,.$
These operators provide a scaled representation of $\mathfrak{so}(3)$, namely
$\mathrm{BT}_{k}(x_{1}),\mathrm{BT}_{k}(x_{2})]=\frac{2i}{k+2}\mathrm{BT}_{k}(x_{3}),$ (13)
with cyclic permutations. This reflects the well-known irreducible representation of $\mathfrak{so}(3)$ on $\mathbb{C}_{k}[z_{1},z_{2}]$, corresponding to the $\mathrm{SU}(2)$ group representation where the action of $A\in\mathrm{SU}(2)$ is $(A\cdot p)(z)=p(A^{-1}z)$. In his original publication, Berezin [4] proves that any quantization on the sphere that is equivariant under the $\mathrm{SO}(3)$ action is unique up to scaling (it is essentially a consequence of the uniqueness of irreducible representations). Thus, we get the connection to the quantization matrix operators $T_{N}(f)$ constructed in section 3.1 above by identifying $\mathbb{C}_{k}[z_{1},z_{2}]\simeq\mathbb{C}^{k+1}$. First, the relation between $N$ and $k$ must be $N=k+1$. Second, it is a matter of getting the scaling right. Recall from section 3.1 that $\left(\sum_{\alpha}T_{N}(x_{\alpha})^{2}\right)^{1/2}=iI$. On the other hand, the operators $\mathrm{BT}_{k}(x_{\alpha})$ fulfill
$\sum_{\alpha=1}^{3}\mathrm{BT}_{k}(x_{\alpha})^{2}=\frac{k}{k+2}.$
Thus, we arrive at the following.
Lemma 20.
For $f\in C^{\infty}(\mathbb{S}^{2})$, let $T_{N}(f)$ denote the matrix quantization operator constructed in section 3.1 above. Then
$T_{N}(f)\simeq i\sqrt{\frac{N+1}{N-1}}\mathrm{BT}_{N-1}(f).$
From (13), the scaling in Lemma 20 recovers that $\{\cdot,\cdot\}$ is approximated, via $T_{N}$, by $\frac{1}{\hbar}[\cdot,\cdot]$ with $\hbar$ as in section 3.1. In particular, since $\sqrt{\frac{N+1}{N-1}}=\mathcal{O}(1)$ as $N\to\infty$, Theorem 3 above by Charles and Polterovich [8] is valid for $T_{N}$, although it was formulated for $\mathrm{BT}_{k}$.